OpenStudy (anonymous):

A customer buys 17.01 in gas and requests one five dollar [$5] lottery ticket, two one dollar [$1] lottery tickets, and one [$3] lottery ticket. He gives you two winning tickets to be redeemed: one for ($5) and the other for ($2). How much change would he receive from a $100 bill? A.$20.1 B.$72.01 C.$74.99 D.$79.99

OpenStudy (anonymous):

17.01+5+2+3=27.01 100-27.01=72.99 72.99+7=79.99

OpenStudy (thesecret20111):

Just write all the information out and solve from there: $100-($17.01+$5+$2+$3)+($5+$2) = $79.99 same as above ^